OpenStudy (anonymous):

how would I find the integral of x^3 dx from 1 to 0?

OpenStudy (anonymous):

anti derivative of \(x^3\) is \(\frac{x^4}{4}\) evaluate at 1 and get \(\frac{1}{4}\)

OpenStudy (anonymous):

ah from one to zero! so it is \(-\frac{1}{4}\) because you are going backwards

OpenStudy (anonymous):



OpenStudy (anonymous):

I need to show that this works for the formula for the sum of n cubes.... just like we did for the sum of squares.....integral of x^2 from 0 to 1 = lim [(n+1)(2n+1)]/6n^2....it says to use this algorithm to evaluate the integral of x^3 from 0 to 1

OpenStudy (anonymous):

the riemann sum?

OpenStudy (anonymous):

think he used that..

OpenStudy (anonymous):

oh really??? have fun!

OpenStudy (anonymous):

we can do it if you like, it will take a while

OpenStudy (anonymous):

first you have to divide the unit interval up in to n pieces, each of which is of length \(\frac{1}{n}\)

OpenStudy (anonymous):

that is your \(\Delta x\)

OpenStudy (anonymous):

he said to "evaluate" and I believe it involves just using the formula we found for the sum of the fist n cubes....

OpenStudy (anonymous):

the we have to say what the \(x_i\) are. you have \(x_0=0, x_1=\frac{1}{n}, x_2=\frac{2}{n}, x_3=\frac{3}{n}\) and in general \(x_i=\frac{i}{n}\)

OpenStudy (anonymous):

then we have to compute \[\sum_{i=0}^nf(x_i)\Delta x\]

OpenStudy (anonymous):

we get \[\sum_{i=1}^n(\frac{i}{n})^3\frac{1}{n}\]